OpenStudy (anonymous):

A book with a mass of 2.0 kg is held in equilibrium on a board with a slope of 60.0° by a horizontal foce. What is the normal force exerted on the book?

OpenStudy (anonymous):

Consider normal force without the slope first. We know this is F=ma. m=2.0 kg a =9.8 m/s (gravity) F = 2*9.8 = 19.6 N So, if the angle is 90 degrees (flat on a table, for example) 19.6 N is the answer. In order to get this in terms of a 60 degree angle, you have to include a geometrical function such as sine or cosine. 19.6 N * sin(60) = 16.974 N We use sine because this is a y-component of the free body diagram, and is not in translation.